1. How much iron chloride (III) is found in 20 ml 4% - solution, density of which 1,133 g/ml?

2. Calculate the molarity of sodium chloride, in 50 ml of which is 2.34 g

dissolved matter.

3. Calculate a 20% molarity - a solution of zinc chloride, the density of which is 1,186 g/cm.

4. What weight of 20% - CON solution should be added to 1 kg 50% solution to

get 25% solution?

5. To 1 litre 20% sodium hydroxide solution, which has a density of 1,225 g/ml, added 10 liters

water. Calculate the percentage concentration of the resulting solution


1
Expert's answer
2021-05-03T07:32:23-0400

1. 20 ml × 0.04

= 0.8 ml

Mass = D × V

= 0.8 × 1133

= 906.4g

2. Molar Mass of NaCl2 = 58.44

= 50/58.44

= 0.86/0.05 + 0.00234

= 16.43 mol/L

3.molar Mass of zinc chloride = 136.286

Density of zinc chloride = 2.91

= 136.286 × 2.91

= 396.59

1186/396.59

= 2.99 × 0.2

= 0.6 g/mol